OpenStudy (anonymous):

six processors are randomly selected from a lot of 100 processors among which 10 are defected.Find the probability of obtaining atleast 1 defective ?

OpenStudy (cwtan):

6 processor slected 10 defective??????

OpenStudy (cwtan):

O.o i get it

OpenStudy (anonymous):

i mean atleast one out of 6 is defective

OpenStudy (cwtan):

P(at least 1 defective)=1-P(NO defective) =1-(0.9)^6

OpenStudy (kropot72):

The hypergeometric distribution applies, the reason being that it is sampling without replacement. P(1 defective) = 0.3687 I will post calculation.

OpenStudy (anonymous):

@cwtan can you explain it more

OpenStudy (kropot72):

\[\frac{\left(\begin{matrix}10 \\ 1\end{matrix}\right)\left(\begin{matrix}90 \\ 5\end{matrix}\right)}{\left(\begin{matrix}100 \\ 6\end{matrix}\right)}\]=0.3687